OpenStudy (callielovesyhuu):

What do the length of the cord and gravity determine for a pendulum? A. velocity B. amplitude C. period D. speed @phi @ganeshie8 @pooja195 @johnweldon1993

OpenStudy (t-dawg02):

\(T=2\pi \sqrt{\frac{L}{g}}\) That equation is the only reason I can answer this equation. If I understood pendulums, I'd help you out with that. There, mathematically, the PERIOD is a function of the *length of the cord* and gravitational acceleration (which depends on *gravity*).

OpenStudy (t-dawg02):

@callielovesyhuu

OpenStudy (callielovesyhuu):

i got the answer

OpenStudy (t-dawg02):

Oh okay